Three people, each working at the same rate, finish a job in 20 hours. At this rate, how many hours will it take five people to finish three such jobs?
Answer by ptaylor(2048)


Let x=amount of time it takes five people to finish one job, then
3x=amount of time it takes five people to finish 3 such jobs
ok. Three people work at the rate of 1/20 of the job per hour, so one person works at the rate of (1/20)/3 or 1/60 of the job per hour
Five people work at the rate of 5*(1/60) or 5/60 of the job per hour
So, our equation to solve is:
(5/60)*x=1 (1 job, that is)
5x=60
x=12 hours--------------amount of time it takes five people to complete 1 job
3x=3*12=36 hours ------------amount of time it takes five people to complete three such jobs
Another way to look at this problem:
If three persons can complete the job in 20 hours, then the job requires 3*20 or 60 person-hours to complete. Three such jobs would require 3*60 or 180 person-hours to complete
So, five persons could complete three such jobs in 180/5=36 hours
Hope this helps----ptaylor